# Localized excitation on the Jacobi elliptic periodic background for the (n+1)-dimensional generalized Kadomtsev-Petviashvili equation

### v1

@@ Abstract
In this paper, the linear spectral problem, which associated with the (n+1)-dimensional generalized Kadomtsev-Petviashvili (gKP) equation, with the Jacobi elliptic function as the external potential is investigated based on the Lamé function, from which some novel local nonlinear wave solutions on the Jacobi elliptic function have been obtained by Darboux transformation, and the corresponding dynamics have also been discussed. At the same time, the degenerate solutions of the nonlinear wave solutions on the Jacobi function background for the gKP equation are constructed by taking the modulus of the Jacobi function to be 0 and 1. The findings indicate that there can be various types of nonlinear wave solutions with different ranges of spectral parameters, including soliton and breather waves. These results will be useful for elucidating and predicting nonlinear phenomena in related physical fields, such as physical ocean and Bose-Einstein condensates.
@@ 1 Introduction
Nonlinear science, renowned for its ability to unravel complex mysteries, is crucial in elucidating a range of scientific phenomena. In recent years, the study of nonlinear waves and their underlying dynamics arose in diverse physical fields, such as fluid mechanics [1], nonlinear optics [2, 3], Bose-Einstein condensate [4], plasma physics [5], deep water [6] and even financial markets[7], have received increasing attentions. Generally, the nonlinear evolution equation can serve as a standard prototype for the characterization of nonlinear waves. Especially, the integrable equation, which can be associated with the linear spectral problem and can be solved by many effective techniques such as Hirota bilinear method [8], inverse scattering transformation method [9, 10, 11], Bäcklund transformation (BT) method [12, 13], Darboux transformation (DT) method [14, 15, 16, 17], algebro-geometric method [18, 19], and the newly favored method of deep learning [20, 21, 22], has always been used as the fundamental governing equation for describing the propagation and dynamics of the nonlinear wave.
@@
However, the majority of current research primarily concentrates on various nonlinear wave solutions on the constant background, such as the extensively studied soliton, periodic wave, breather and rogue wave solutions are all based on zero and plane wave backgrounds. It is quite natural that nonlinear waves on the nonconstant wave background and their corresponding underlying dynamics should also be discussed, which is also the main motivation of our paper. Of course, there have been many scholars who have paid attention to this issue. Recently, Kedziora and Chin et al. investigated nonlinear wave solutions on the periodic wave backgrounds based on DT and numerical method [23, 24]; Lou and Lin derived the interaction solutions between soliton and periodic wave solutions by nonlocal symmetry and dressing method [25, 26]; Chen and Pelinovsky et al. constructed many nonlinear wave solutions on the Jacobi elliptic function periodic wave backgrounds for some classical Nonlinear Schrödinger (NLS) type equations [27, 28, 29, 30, 31, 32, 33] by the nonlinearization technique [34] and DT; Feng and Ling et al. discussed multi-breather and high-order rogue waves on the elliptic function background by DT method and theta functions [35]; Using the DT and Baker-Akhiezer function method, Li and Geng constructed explicit solutions on the periodic backgrounds [36, 37, 38, 39]; Hoefer et al. gave the bright and dark breather solutions on the cnoidal wave background by DT method [40]; Chen et al. considered the rogue wave on the periodic background by PINN deep learning method [41]; We also discussed some nonlinear wave solutions on the periodic background by DT method [42, 43, 44].
@@
For several decades, physics has been concerned with nonlinear local excitations on the nonconstant background, especially periodic wave background. In fluid mechanics, a new type of solitary wave motion in incompressible fluids of non-uniform density has been studied by Davis and Acrivos in both experiments and theory [45]. Following this, in the study of internal waves, Farmer and Smith observed solitary waves followed by a sequence of periodic waves [46]. Xu and Chabchoub et al. studied the modulation instability and associated rogue breathers on periodic wave background experimentally [47]. At the same time, the related nonlinear wave theories and laboratory experiments have shown that nonconstant background can better describe physical reality, in which a very typical background is periodic wave, initially appeared as solutions to the conservation laws governing shallow water and subsequently manifested in the domains of ion plasmas [48], nonlinear optics [49], and elastic solids [50]. These mathematic-physical results lead to an increased interest in the study of nonlinear wave solutions on the periodic wave background.
@@
To describe more complex and rich physical phenomena in real physical situations or experiments, high-dimensional integrable systems are typically used as prototypes, which possess richer mathematical structures and dynamic properties, although they are difficult to handle [51, 52]. Therefore, it is a very natural and meaningful problem to study nonlinear waves on the periodic wave background for the high-dimensional integrable systems with physical significance. Under this motivation, this paper concerns the (n+1)-dimensions generalized Kadomtsev-Petviashvili (gKP) equation [53]:
@@
$\quad+\frac{\gamma}{\delta}\partial_{x_{1}}{u_{x_{2}}}-\sum_{i=2}^{n}\sigma_{i}\partial_{x_{i}},$ (3b)
@@
can be derived from equation (1) under the parameters $\beta=1,\,\sigma_{i}=0,\,\gamma=\pm 3$, and the (3+1)-dimensional generalized KP equation[54]
@@
Previous studies of the nonlinear waves on the periodic wave background for the KP family equation mainly focus on Weierstrass elliptic function wave. Krichever investigated the theory of elliptic solitons for the KP equation and the corresponding Calogero-Moser system [56]; Brezhnev studied the solutions for Lamé equation (6) by Hermite Ansatz method [57]; Zhang and Li proposed the bilinear framework for elliptic soliton solutions, which are composed of the Lamé-type plane wave factors [58]. Nijhoff, Sun, and Zhang established an infinite family of solutions in terms of elliptic functions of the lattice Boussinesq systems [59]; However, to the best of our knowledge, the nonlinear waves on the Jacobi elliptic function periodic wave background for the (n+1)-dimensional gKP equation (1) remains unexplored, under which the linear spectral problem (2) can be connected with the Lamé equation in Jacobi form [55],
@@
This paper is structured as follows. In Sec. 2, the $N$-th DT for the (n+1)-dimensional gKP equation (1) is constructed directly. In Sec. 3, the Jacobi elliptic periodic seed solution of the gKP equation (1) is given using the Jacobi elliptic function expansion method. In Sec. 4, a general solution of the linear spectral problem with Jacobi function coefficients is constructed, and the spectrum is analyzed. In Sec. 5, various nonlinear wave solutions on the Jacobi elliptic periodic function for the (n+1)-dimensional gKP equation (1) are derived via one-time and two-times DT. In Sec. 6, degenerate solutions have been studied by modulus $k$ tending to 00 or $1$, where soliton solutions and interaction solutions can be obtained. Finally, conclusions and future work are summarized in Sec. 7.
@@ 2 Lax pair and Darboux transformation for the gKP equation
$\varphi[1]_{t}=\left(\frac{\gamma}{\delta}+\beta\delta\right)\varphi[1]_{x_{1}x_{2}}+\left(\frac{\gamma}{\delta}\partial_{x_{1}}{u[1]_{x_{2}}}\right.$
@@
$\varphi_{j}[l]_{t}=\left(\frac{\gamma}{\delta}+\beta\delta\right)\varphi_{j}[l]_{x_{1}x_{2}}+\left(\frac{\gamma}{\delta}\partial_{x_{1}}{u[l]_{x_{2}}}\right.$
@@
$\varphi[N]_{t}=\left(\frac{\gamma}{\delta}+\beta\delta\right)\varphi[N]_{x_{1}x_{2}}+\left(\frac{\gamma}{\delta}\partial_{x_{1}}{u[N]_{x_{2}}}\right.$
@@
$=u[N-2]+2\left(\ln\mathrm{Wr}\left[\varphi_{N-1}[N-2],\right.\right.$
@@ 3 Traveling Jacobi elliptic function seed solution for the gKP equation
By applying the Jacobi elliptic function expansion method [60, 61], $U(\eta)$ can be expressed as a finite series of Jacobi elliptic functions of the form:
@@ 4 Solutions of the spectral problems via Lamé equation
where $\nu=-\lambda$. Obviously, equation (20a) can be associated with the Lamé equation (6). According to the general solution of the Lamé equation [55], two linearly independent solutions of system (20) can be assumed as
@@
$\varphi_{1\pm}(\eta):=\mathrm{dn}(\eta,k)\mp\omega_{2}y,\quad\lambda_{1\pm}=-k^{2}\mp\delta\omega_{2},$ (24)
$\varphi_{2\pm}(\eta):=\mathrm{cn}(\eta,k)\mp\omega_{2}y,\quad\lambda_{2\pm}=-1\mp\delta\omega_{2},$
$\varphi_{3\pm}(\eta):=\mathrm{sn}(\eta,k)\mp\omega_{2}y,\quad\lambda_{3\pm}=-1-k^{2}\mp\delta\omega_{2},$
@@
Therefore, $\lambda_{+}$ and $\lambda_{-}$ can be divided into four intervals according to the relations (25). Based on the characteristic equation (23), the definition of $\alpha$ can be given in each interval when $\lambda_{+}$ is increased from $\lambda_{3+}$ to +$\infty$+ ∞.
@@
Employing the properties of Jacobi elliptic functions [62] yields
@@
Using the special relations of Jacobi elliptic functions [62]
@@
×(±$\quad\times\left(\pm\frac{H^{\prime}(\eta\pm\alpha)}{H(\eta\pm\alpha)}\mp Z(\eta)-Z(\alpha)\right)-\sum_{i=2}^{n}\sigma_{i}\omega_{i}$
±$\quad\pm 4k^{2}\mathrm{sn}(\eta,k)\mathrm{cn}(\eta,k)\mathrm{dn}(\eta,k),$
@@
Using properties of the Jacobi theta function, the case of $\omega_{1}$ in the three different cases (i), (ii), (iii) ca be given in the following proposition.
@@
The half-periodic transformation of the Jacobi theta function and the logarithmic derivative formula are mainly utilized [63], and the proof procedure is not repeated here. Up to this point, the general solution of system (20) is a linear combination of two linearly independent solutions (21), we have the following theorem.
@@
5 Breather waves on the traveling Jacobi elliptic function periodic wave background
@@ 5.1 First order breather waves on the traveling Jacobi elliptic function periodic wave background
$-k^{2}\mathrm{sn}^{2}(\eta,k)=\frac{E(k)}{K(k)}-1+\partial_{x_{1}}^{2}\ln\Theta(\eta).$ (38)
@@
From the properties of the Jacobian function [62] it can be deduced that
@@
$-k^{2}\mathrm{sn}^{2}(\eta,k)=-\frac{\Theta^{\prime\prime}(0)}{\Theta(0)}+\partial_{x_{1}}^{2}\ln\Theta(\eta).$ (41)
@@
Taking $\alpha=K(k)+iK^{\prime}(k)+\gamma$ and $\gamma=F(\phi_{\gamma},k)\in[0,K(k))$ in the general solution (36), and utilizing the half-periodic translations of Jacobi theta functions [64]:
@@
×$\qquad\qquad\quad\times\Theta(\eta+K(k)+\gamma),$ (45b)
@@
×$\qquad\qquad\quad\times\Theta(\eta+K(k)-\gamma).$ (45c)
@@
Furthermore, it can be deduced from the properties [65]
@@
$\kappa_{1}=\mu_{1}\big{(}x_{1}+c_{1}t+\sum_{i=2}^{n}p_{i}x_{i}\big{)},$
$\mu_{1}=\frac{\sqrt{\lambda+1+\delta\omega_{2}+k^{2}}\sqrt{\lambda+\delta\omega_{2}+k^{2}}}{\sqrt{\lambda+1+\delta\omega_{2}}}$
$\quad\ \ \ -Z(\phi_{\gamma},k),$
@@
×$\quad\ \ \ \times\sqrt{\lambda+1+\delta\omega_{2}}-\sum_{i=2}^{n}\sigma_{i}\omega_{i}\big{)}/\mu_{1},$
@@
for $i=2,3,\cdots,n$. $\beta,\delta,\omega_{i},\sigma_{i}\in\mathbb{R}$ are arbitrary, and $\phi_{\gamma}\in[0,\frac{\pi}{2})$. The appropriate normalized phase shift for the background Jacobi elliptic wave is expressed as
@@
$F_{1}=\left(Z_{x_{1}}(\eta-\gamma)-Z_{x_{1}}(\eta+\gamma)\right)\Theta^{2}(\eta-\gamma)$
×$\qquad\ \times\mathrm{exp}(-2\kappa_{1})+\big{(}(Z(\eta+\gamma)-Z(\eta-\gamma)+2\mu_{1})^{2}$
$\qquad\ +Z_{x_{1}}(\eta-\gamma)-Z_{x_{1}}(\eta+\gamma)\big{)}\Theta(\eta+\gamma)\Theta(\eta-\gamma),$
@@
The propagation of the breather wave (53) over different spatial variables $x_{1}$, $x_{2}$, and $x_{3}$ with respect to the time variable $t$ under some given parameter values and the corresponding contour plots are shown in Figure 1. Figure 1(a) demonstrates the bright periodic wave morphology presented on the periodic background wave, which propagates with a velocity of $c$ along the $x_{1}$-$t$ direction, as well as the breather presenting an inverse width $\mu_{1}$ and a velocity $c_{1}$ propagation, and the bright breather is generated with a phase shift of $-2\gamma$. Similarly, in both Figure 1(b) and Figure 1(c), the periodic background wave propagates along the direction of $t$. The inverse widths of the bright breathers generated above are $\omega_{i}(i=2,~{}3)$, the propagation speed are $c_{1}/p_{i}~{}(i=2,~{}3)$ respectively, and the phase shift are both $-2\gamma$. In the same way, the evolution behavior of breather wave in other spatial variable $x_{i}$ and time $t$ can be obtained by taking corresponding parameters in solution (53). These parameters represent the characteristics of the breather wave and provide distinct physical meanings in different physical scenarios.
(a)                                              (b)                                              (c)
Figure 1: Propagation and contour plots for the breather wave (53), where $k=0.3$, $\lambda=1.8$, $\delta=1$, $\omega_{2}=\omega_{3}=1$, $\omega_{i}(i=4,\cdots,n)=0$. (a) $x_{2}=x_{3}=1$; (b) $x_{1}=x_{3}=1$; (c) $x_{1}=x_{2}=1$.
(a)                                              (b)                                              (c)
Figure 2: Propagation and contour plots for nonlinear wave solution (62), where $k=0.6$, $\lambda=-2.1$, $\delta=1$, $\omega_{2}=\omega_{3}=1$, $\omega_{i}(i=4,\cdots,n)=0$. (a) $x_{2}=x_{3}=1$; (b) $x_{1}=x_{3}=1$; (c) $x_{1}=x_{2}=1$.
@@
×$\qquad\qquad\quad\times\Theta(x+\alpha-iK^{\prime}(k)),$ (54a)
@@
×$\qquad\qquad\quad\times\Theta(x-\alpha-iK^{\prime}(k)).$ (54b)
@@
On the other hand, it can be directly deduced from the property [65]
@@
×$\quad\ \ \ \times\sqrt{\lambda+1+\delta\omega_{2}}-\sum_{i=2}^{n}\sigma_{i}\omega_{i}\big{)}/\mu_{1},$

### v2

@@ Abstract
In this paper, the linear spectral problem, which associated with the (n+1)-dimensional generalized Kadomtsev-Petviashvili (gKP) equation, with the Jacobi elliptic function as the external potential is investigated based on the Lamé function, from which some novel local nonlinear wave solutions on the Jacobi elliptic function have been obtained by Darboux transformation, and the corresponding dynamics have also been discussed. The degenerate solutions of the nonlinear wave solutions on the Jacobi function background for the gKP equation are constructed by taking the modulus of the Jacobi function to be 0 and 1. The findings indicate that there can be various types of nonlinear wave solutions with different ranges of spectral parameters, including soliton and breather waves. Furthermore, the interplay between nonlinearity and dispersion is found to have observable effects on the propagation dynamics of breather waves. These results will be useful for elucidating and predicting nonlinear phenomena in related physical fields, such as fluid mechanics and physical ocean.
@@ 1 Introduction
Nonlinear science, renowned for its ability to unravel complex mysteries, is crucial in elucidating a range of scientific phenomena. In recent years, the study of nonlinear waves and their underlying dynamics arose in diverse physical fields, such as fluid mechanics AD-1988 , nonlinear optics AH-2002 , BA-2015 , Bose-Einstein condensate LP-SS-2016 , plasma physics HB-2011 , deep water PAM-HBB-2002 and even financial marketsZY-2011 , have received increasing attentions. Generally, the nonlinear evolution equation can serve as a standard prototype for the characterization of nonlinear waves. Especially, the integrable equation, which can be associated with the linear spectral problem and can be solved by many effective techniques such as Hirota bilinear method RH-2004 , inverse scattering transformation method MB-FP-2002 , AG-AD-2014 , GQZ-ZYY-2020 , Bäcklund transformation (BT) method CR-WK-2002 , JMT-SFT-MJX-2016 , Darboux transformation (DT) method MVB-SMA-1991 , XYW-YQY-2015 , XYW-ZY-2020 , CLT-ZWW-2021 , algebro-geometric method FG-2003 , FG-2008 , and the newly favored method of deep learning JCP-YC-2023 , SNL-YC-2022 , JHL-JCC-2022 , has always been used as the fundamental governing equation for describing the propagation and dynamics of the nonlinear wave.
@@
However, the majority of current research primarily concentrates on various nonlinear wave solutions on the constant background, such as the extensively studied soliton, periodic wave, breather and rogue wave solutions are all based on zero and plane wave backgrounds. It is quite natural that nonlinear waves on the nonconstant wave background and their corresponding underlying dynamics should also be discussed, which is also the main motivation of our paper. Of course, there have been many scholars who have paid attention to this issue. Recently, Kedziora and Chin et al. investigated nonlinear wave solutions on the periodic wave backgrounds based on DT and numerical method DJK-AA-2014 , SAC-OAA-2017 ; Lou and Lin derived the interaction solutions between soliton and periodic wave solutions by nonlocal symmetry and dressing method SYL-XPC-2014 , JL-XWJ-2018 ; Chen and Pelinovsky et al. constructed many nonlinear wave solutions on the Jacobi elliptic function periodic wave backgrounds for some classical Nonlinear Schrödinger (NLS) type equations JBC-DE-2018 , JBC-DE-2019 , JBC-RSZ-2020 , JBC-DEP-2021 , JBC-DEP-REW-2020 , HQZ-FC-ZJP-2021 , HQZ-FC-2021 by the nonlinearization technique CWC-XGG-1990 and DT; Feng and Ling et al. discussed multi-breather and high-order rogue waves on the elliptic function background by DT method and theta functions BFF-LML-DAT-2020 ; Using the DT and Baker-Akhiezer function method, Li and Geng constructed explicit solutions on the periodic backgrounds RML-XGG-2020 , RL-2023 , XGG-RML-2022 , RML-JRG-XGG-2023 ; Hoefer et al. gave the bright and dark breather solutions on the cnoidal wave background by DT method MAH-AM-2023 ; Chen et al. considered the rogue wave on the periodic background by PINN deep learning method WQP-JCP-2022 ; We also discussed some nonlinear wave solutions on the periodic background by DT method YQY-HHD-2021 , JBL-YQY-2024 , XMX-YQY-2024 .
@@
For several decades, physics has been concerned with nonlinear local excitations on the nonconstant background, especially periodic wave background. In fluid mechanics, a new type of solitary wave motion in incompressible fluids of non-uniform density has been studied by Davis and Acrivos in both experiments and theory RED-AA-1967 . Following this, in the study of internal waves, Farmer and Smith observed solitary waves followed by a sequence of periodic waves DMF-JDS-1980 . Xu and Chabchoub et al. studied the modulation instability and associated rogue breathers on periodic wave background experimentally GX-AC-2020 . At the same time, the related nonlinear wave theories and laboratory experiments have shown that nonconstant background can better describe physical reality, in which a very typical background is periodic wave, initially appeared as solutions to the conservation laws governing shallow water and subsequently manifested in the domains of ion plasmas JSB-VN-1994 , nonlinear optics JLS-GJS-1997 , and elastic solids CM-JS-JR-2019 . These mathematic-physical results lead to an increased interest in the study of nonlinear wave solutions on the periodic wave background.
@@
To describe more complex and rich physical phenomena in real physical situations or experiments, high-dimensional integrable systems are typically used as prototypes, which possess richer mathematical structures and dynamic properties, although they are difficult to handle CLT-ZWW-2019 , YQY-XW-2015 . Therefore, it is a very natural and meaningful problem to study nonlinear waves on the periodic wave background for the high-dimensional integrable systems with physical significance. Under this motivation, this paper concerns the (n+1)-dimensions generalized Kadomtsev-Petviashvili (gKP) equation GQX-AMW-2023 :
@@
$\quad+\frac{\gamma}{\delta}\partial^{-1}_{x_{1}}{u_{x_{2}}}-\sum_{i=2}^{n}\sigma_{i}\partial_{x_{i}},$ (3b)
@@
can be derived from equation (1) under the parameters $\beta=1,\,\sigma_{i}=0,\,\gamma=\pm 3$, and the (3+1)-dimensional generalized KP equationLLH-YFY-YC
@@
Previous studies of the nonlinear waves on the periodic wave background for the KP family equation mainly focus on Weierstrass elliptic function wave. Krichever investigated the theory of elliptic solitons for the KP equation and the corresponding Calogero-Moser system IMK-1994 ; Brezhnev studied the solutions for Lamé equation (6) by Hermite Ansatz method YVB-2004 ; Zhang and Li proposed the bilinear framework for elliptic soliton solutions, which are composed of the Lamé-type plane wave factors XL-DJZ-2020 . Nijhoff, Sun, and Zhang established an infinite family of solutions in terms of elliptic functions of the lattice Boussinesq systems FWN-YYS-2023 ; However, to the best of our knowledge, the nonlinear waves on the Jacobi elliptic function periodic wave background for the (n+1)-dimensional gKP equation (1) remains unexplored, under which the linear spectral problem (2) can be connected with the Lamé equation in Jacobi form ELI-1956 ,
@@
This paper is structured as follows. In Sec. 2, the $N$-th DT for the (n+1)-dimensional gKP equation (1) is constructed directly. In Sec. 3, the Jacobi elliptic periodic seed solution of the gKP equation (1) is given using the Jacobi elliptic function expansion method. In Sec. 4, a general solution of the linear spectral problem with Jacobi function coefficients is constructed, and the spectrum is analyzed. In Sec. 5, various nonlinear wave solutions on the Jacobi elliptic periodic function for the (n+1)-dimensional gKP equation (1) are derived via one-time and two-times DT. In Sec. 6, degenerate solutions have been studied by modulus $k$ tending to 0 or $1$, where soliton solutions and interaction solutions can be obtained. Finally, conclusions and future work are summarized in Sec. 7.
@@ 2 Lax pair and Darboux transformation for the gKP equation
$\varphi[1]_{t}=\left(\frac{\gamma}{\delta}+\beta\delta\right)\varphi[1]_{x_{1}x_{2}}+\left(\frac{\gamma}{\delta}\partial^{-1}_{x_{1}}{u[1]_{x_{2}}}\right$.
@@
$\varphi_{j}[l]_{t}=\left(\frac{\gamma}{\delta}+\beta\delta\right)\varphi_{j}[l]_{x_{1}x_{2}}+\left(\frac{\gamma}{\delta}\partial^{-1}_{x_{1}}{u[l]_{x_{2}}}\right$.
@@
$\varphi[N]_{t}=\left(\frac{\gamma}{\delta}+\beta\delta\right)\varphi[N]_{x_{1}x_{2}}+\left(\frac{\gamma}{\delta}\partial^{-1}_{x_{1}}{u[N]_{x_{2}}}\right$.
@@
[…]
@@ 3 Traveling Jacobi elliptic function seed solution for the gKP equation
By applying the Jacobi elliptic function expansion method EGF-BYH-2002 , ZYY-2003 , $U(\eta)$ can be expressed as a finite series of Jacobi elliptic functions of the form:
@@ 4 Solutions of the spectral problems via Lamé equation
where $\nu=-\lambda$. Obviously, equation (20a) can be associated with the Lamé equation (6). According to the general solution of the Lamé equation ELI-1956 , two linearly independent solutions of system (20) can be assumed as
@@
$\varphi_{1\pm}(\eta)=\mathrm{dn}(\eta,k)\mp\omega_{2}y,\quad\lambda_{1\pm}=-k^{2}\mp\delta\omega_{2},$ (24)
$\varphi_{2\pm}(\eta)=\mathrm{cn}(\eta,k)\mp\omega_{2}y,\quad\lambda_{2\pm}=-1\mp\delta\omega_{2},$
$\varphi_{3\pm}(\eta)=\mathrm{sn}(\eta,k)\mp\omega_{2}y,\quad\lambda_{3\pm}=-1-k^{2}\mp\delta\omega_{2},$
@@
Therefore, $\lambda_{+}$ and $\lambda_{-}$ can be divided into four intervals according to the relations (25). Based on the characteristic equation (23), the definition of $\alpha$ can be given in each interval when $\lambda_{+}$ is increased from $\lambda_{3+}$ to $\infty$.
@@
Employing the properties of Jacobi elliptic functions DFL-2013 yields
@@
Using the special relations of Jacobi elliptic functions DFL-2013
@@
$\quad\times\left(\pm\frac{H^{\prime}(\eta\pm\alpha)}{H(\eta\pm\alpha)}\mp Z(\eta)-Z(\alpha)\right)-\sum_{i=2}^{n}\sigma_{i}\omega_{i}$
$\quad\pm 4k^{2}\mathrm{sn}(\eta,k)\mathrm{cn}(\eta,k)\mathrm{dn}(\eta,k),$
@@
Using properties of the Jacobi theta function, the case of $\omega_{1}$ in the three different cases (i), (ii), (iii) can be given in the following proposition.
@@
The half-periodic transformation of the Jacobi theta function and the logarithmic derivative formula are mainly utilized JB-CS-2019 , and the proof procedure is not repeated here. Up to this point, the general solution of system (20) is a linear combination of two linearly independent solutions (21), we have the following theorem.
@@
5 Breather Waves on the Traveling Jacobi Elliptic Function Periodic Background with Dispersion Effects
@@ 5.1 First order breather waves on the traveling Jacobi elliptic function periodic wave background
$k^{2}\mathrm{sn}^{2}(\eta,k)=\frac{E(k)}{K(k)}-1+\partial_{x_{1}}^{2}\ln\Theta(\eta).$ (38)
@@
From the properties of the Jacobian function DFL-2013 it can be deduced that
@@
$k^{2}\mathrm{sn}^{2}(\eta,k)=-\frac{\Theta^{\prime\prime}(0)}{\Theta(0)}+\partial_{x_{1}}^{2}\ln\Theta(\eta).$ (41)
@@
Taking $\alpha=K(k)+iK^{\prime}(k)+\gamma$ and $\gamma=F(\phi_{\gamma},k)\in[0,K(k))$ in the general solution (36), and utilizing the half-periodic translations of Jacobi theta functions ISG-IMR-2014 :
@@
$\qquad\qquad\quad\times\Theta(\eta+K(k)+\gamma),$ (45b)
@@
$\qquad\qquad\quad\times\Theta(\eta+K(k)-\gamma).$ (45c)
@@
Furthermore, it can be deduced from the properties PFB-MDF-2013
@@
$\kappa_{1}=\mu_{1}\big(x_{1}+c_{1}t+\sum_{i=2}^{n}p_{i}x_{i}\big),\qquad D=-\sum_{i=2}^{n}\sigma_{i}\omega_{i}$
$\mu_{1}=\frac{\sqrt{\lambda+1+\delta\omega_{2}+k^{2}}\sqrt{\lambda+\delta\omega_{2}+k^{2}}}{\sqrt{\lambda+1+\delta\omega_{2}}}-Z(\phi_{\gamma},k),$
@@
$\quad\ \ \ \times\sqrt{\lambda+1+\delta\omega_{2}}+D\big)/\mu_{1},$
@@
for $i=2,3,\cdots,n$. $\beta,\delta,\omega_{i},\sigma_{i}\in\mathbb{R}$ are arbitrary, $D$ is the dispersion parameter and $\phi_{\gamma}\in[0,\frac{\pi}{2})$. The appropriate normalized phase shift for the background Jacobi elliptic wave is expressed as
@@
$F_{1}=\left(Z_{x_{1}}(\eta-\gamma)-Z_{x_{1}}(\eta+\gamma)\right)\Theta^{2}(\eta-\gamma)\mathrm{exp}(-2\kappa_{1})$
$\qquad\ +\big((Z(\eta+\gamma)-Z(\eta-\gamma)+2\mu_{1})^{2}+Z_{x_{1}}(\eta-\gamma)$
$\qquad\ -Z_{x_{1}}(\eta+\gamma)\big)\Theta(\eta+\gamma)\Theta(\eta-\gamma),$
@@
It is worth noting that the explicit breather solution given in Theorem 4 contains rich physical information through its parameter dependence. In particular, although the additional linear dispersion terms in Eq. (1), namely $\sum_{i=2}^{n}\sigma_{i}u_{x_{1}x_{i}}$, are algebraically simple and can be removed by a Galilean-type coordinate transformation, they enter the solution explicitly through the breather velocity. For the bright breather solution (53), the propagation velocity $c_{1}$ appearing in (51) depends linearly on the effective dispersion parameter $D=-\sum_{i=2}^{n}\sigma_{i}\omega_{i}$, which implies that the longitudinal propagation of the breather can be continuously tuned by adjusting the linear dispersion coefficients. This velocity modulation reflects the coupling between longitudinal and transverse dispersive channels in the underlying spectral structure, while the intrinsic waveform profile of the breather remains unchanged.
The propagation of the breather wave (53) with respect to the time variable $t$ along different spatial directions $x_{1}$, $x_{2}$, and $x_{3}$, together with the corresponding contour plots, is illustrated in Fig. 1 for a representative set of parameters. Figure 1(a) shows a bright breather embedded in a periodic background wave, propagating with velocity $c$ along the $x_{1}$$t$ plane. The breather itself is characterized by an inverse width $\mu_{1}$, a propagation velocity $c_{1}$, and a phase shift of $2\gamma$. Similarly, Figs. 1(b) and 1(c) display the evolution of the same solution in the $x_{2}$$t$ and $x_{3}$$t$ planes, respectively. In these cases, the inverse widths of the bright breathers are given by $\omega_{i}~(i=2,3)$, while their propagation velocities are $c_{1}/p_{i}~(i=2,3)$, both accompanied by the same phase shift $2\gamma$. By appropriately choosing the parameters in solution (53), the dynamical behaviors of the breather wave along other spatial directions can be readily obtained, providing flexible physical interpretations in different multidimensional settings.
Figure 1: Propagation profiles and corresponding contour plots of the breather wave solution (53). The parameters are chosen as $k=0.3$, $\alpha=\pi/3$, $\beta=1$, $\lambda=1.8$, $\delta=1$, $\omega_{2}=\omega_{3}=1$, and $D=5$. (a) Evolution in the $x_{1}$$t$ plane with $x_{2}=x_{3}=1$; (b) evolution in the $x_{2}$$t$ plane with $x_{1}=x_{3}=1$; (c) evolution in the $x_{3}$$t$ plane with $x_{1}=x_{2}=1$.
Figure 2: Propagation profiles and corresponding contour plots of the dark breather solution (62). The parameters are chosen as $k=0.6$, $\alpha=\pi/3$, $\beta=1$, $\lambda=-2.1$, $\delta=1$, $\omega_{2}=\omega_{3}=1$, and $D=5$. (a) Evolution in the $x_{1}$$t$ plane with $x_{2}=x_{3}=1$; (b) evolution in the $x_{2}$$t$ plane with $x_{1}=x_{3}=1$; (c) evolution in the $x_{3}$$t$ plane with $x_{1}=x_{2}=1$.
@@
$\qquad\qquad\quad\times\Theta(x+\alpha-iK^{\prime}(k)),$ (54a)
@@
$\qquad\qquad\quad\times\Theta(x-\alpha-iK^{\prime}(k)).$ (54b)
@@
On the other hand, it can be directly deduced from the property PFB-MDF-2013
@@
$\quad\ \ \ \times\sqrt{\lambda+1+\delta\omega_{2}}+D)/\mu_{2},$